how many liters of water must be added to 20 liters of a 70% antifreeze solutionto produce a 50% solution?: how many liters of water must be added to 20 liters of a 70% antifreeze solutionto produce a 50% solution?
Answer by stanbon(18327)

how many liters of water must be added to 20 liters of a 70% antifreeze solution to produce a 50% solution?
------------------
EQUATION:
water + water = water
(0.30)*20 + x = (0.50)(20+x)
6 + x = 10 + (1/2)x
(1/2)x = 4
x = 8 liters of water must be added
=================
Cheers,
Stan H.